a number consist of two digits whode sum is 15 . if 9 is added to the number then the digits change places find the number​

Grade:11

1 Answers

Himanshu saxena
2 years ago
Lets, say the original number has “x” at the 10th place and “y” at the one’s place,
thus the number can be written as (10x + y)
also, (x + y = 15; given) …... Eq(i)
also, (10x + y +9 = 10y + x) ….. Eq(ii)
Solving Eq(ii) gives us,
(x – y = -1) ...... Eq(iii)
Adding, Eq(i) and Eq(iii), we get x= 7 thus, y=8
Hence, the original number is (10 * 7 + 8 = 78).
 
The original number is 78.